Provide a structure for the following compound: C10H13NO2; IR: 3285, 1659, 1246 cm?1; 1H NMR spectrum:
Answer #1
Concepts and reason

The concept is to identify the structure of the compound from given 1HNMR^{\rm{1}}{\rm{HNMR}} spectrum.

Fundamentals

Number of signals in the 1HNMR^{\rm{1}}{\rm{HNMR}} spectrum is equal to number of non-equivalent hydrogens. If proton attached to more electronegative atom (i.e. O, N, X), then it gives signal at more downfield from the reference compound (i.e. TMS). IR spectrum is useful to identify the functional groups.

Calculate the double bond equivalence is as follows.

DBE=NC+1((NH+NCl)NN2)ThemolecularformulaforthegivencompoundisC10H13NO2DBE=10+1((13+0)12)DBE=5So,thecompoundhas5doublebonds,itindicatesthepresenceofbenzenering.\begin{array}{l}\\{\rm{DBE}}\,{\rm{ = }}\,{{\rm{N}}_{\rm{C}}}\,{\rm{ + }}\,{\rm{1}}\, - \left( {\frac{{{\rm{(}}{{\rm{N}}_{\rm{H}}}\,{\rm{ + }}\,{{\rm{N}}_{{\rm{Cl}}}}{\rm{)}} - {{\rm{N}}_{\rm{N}}}}}{{\rm{2}}}} \right)\\\\{\rm{The}}\,{\rm{molecular}}\,{\rm{formula}}\,{\rm{for}}\,{\rm{the}}\,{\rm{given}}\,{\rm{compound}}\,{\rm{is}}\,\,{{\rm{C}}_{10}}{{\rm{H}}_{{\rm{13}}}}{\rm{N}}{{\rm{O}}_2}\\\\{\rm{DBE}}\,{\rm{ = }}\,10{\rm{ + 1}}\, - \left( {\frac{{{\rm{(13}}\,{\rm{ + }}\,{\rm{0)}} - 1}}{{\rm{2}}}} \right)\\\\{\rm{DBE}}\,{\rm{ = }}\,\,5\\\\{\rm{So,}}\,{\rm{the}}\,{\rm{compound}}\,{\rm{has}}\,\,5\,{\rm{double}}\,{\rm{bonds,}}\,{\rm{it}}\,{\rm{indicates}}\,{\rm{the}}\,{\rm{presence}}\,{\rm{of}}\,{\rm{benzene}}\,{\rm{ring}}.\\\end{array}

Total 6 signals are observed in the given 1HNMR^{\rm{1}}{\rm{HNMR}} spectrum.

Two 2H doublets around 6.9 to 7.5 ppm corresponds a disubstituted benzene ring.

2H quartet at 4.0 ppm and 3H triplet at 1.3 ppm are adjacent to each other.

3H singlet at 2.1 ppm means that the methyl group attached to carbonyl group.

Thus, the structure of unknown compound is,

Ans:

Thus, the structure of unknown compound is,